OpenStudy (anonymous):

| 4x | = |15+x| | 5 | |2y - 1| Select one: a. | 3 4| b. |-3 -4| c. |5 3| d. | -2 5|

OpenStudy (anonymous):

|■(4x@5)|=|■(15+x@2y-1)| 4x=15+x 4x-x=15 X=15/3=5 5=2y-1 2y=5+1 Y=6/2=3

OpenStudy (anonymous):

That's not an answer choice though.

OpenStudy (anonymous):

Nevermind thanks you was right